The height of a projectile fired vertically nto the air (neglecting air resistance) at an initial velocity of 96 feet per second is a function of time, x, and is given by the equation f(x)=96-16x^2. Find the highest point reached by the projectile.

Answer by stanbon(48568)

The height of a projectile fired vertically nto the air (neglecting air resistance) at an initial velocity of 96 feet per second is a function of time, x, and is given by the equation f(x)=96x-16x^2. Find the highest point reached by the projectile.
-----------------------
That is a quadratic equation with a = -16 and b = 96
-----------------
The maximum height occurs when x = -b/(2a) = -96/(-32) = 3 seconds.
-------------------
The height at that time is f(3) = 96*3-16(3)^2 = 144 ft.
============================
Cheers,
Stan H.